OpenStudy (anonymous):

Find each percent increase or decrease to the nearest percent. 1. from 16 to 20 A. Increase B. Decrease what is the nearest percentage ---------------------------------------------------------- 2. from 30 to 24 A. Increase B. Decrease what is the nearest percentage ---------------------------------------------------------- etc ...

OpenStudy (ganpat):

1st one.. what do u think increased or decreased ??

OpenStudy (anonymous):

idk

OpenStudy (ganpat):

look at the values .. 16 tooooooooooo 20.. is it decreasing or increasing ??

OpenStudy (anonymous):

increasing

OpenStudy (ganpat):

so select option A, percentage increase = { (20-16)/16 } * 100 = (4/16) * 100 = 0.25 * 100 = 25%..

OpenStudy (ganpat):

SO now u can do for rest... right ??

OpenStudy (anonymous):

yep thank you :D